Xenon hexafluoride was one of the first noble gas compound synthesized. The solid reacts rapidly with the silicone dioxide and glass or quartz containers to form liquid XeOF4 and gaseous silicone tetrafluoride. What is the pressure in a 1.00 L container at 25°C after 1.86 g xenon hexafluoride react? (Assume that silicone tetrafluoride is the only gas present and that it occupies the entire volume)
Answer #1

Reaction is

SiO2 + 2 XeF6 ....................................> 2 XeOF4 + SiF4

In this reaction,

one mole SiO2 react with 2 mole XeF6 to form 2 mole XeOF4 and one mole SiF4.

thus

1.86 g xenon hexafluoride = mass / molar mass = 1.86 g /  245.28 g/mole = 7.583 * 10^-3 mole.

thus

mole of SiF4 produce = 7.583 * 10^-3 mole / 2 = 3.79 * 10^-3 mole.

we have

V = 1.00 L

T = 298 K

mole of gas (n) = 3.79 * 10^-3 mole.

Ideal gas equation is

PV = nRT

or

P = nRT / V = 3.79 * 10^-3 * 0.0821 * 298 / 1.00 = 0.0927 atm

thus pressure is 0.0927 atm